OpenStudy (anonymous):

Perform a sequence of transformations to obtain the graph of each of the following functions. Do this step by step and explain each step along the way. 2. y = 5x − 10 3. y = √3x + 9 + 5 4. y = 8− √10 − x 5. y = 2(x + 4)^3 6. y = 5 over x (5/x) 7. y = −4 |−2x + 3

OpenStudy (mathteacher1729):

So ... for # 2. Does this mean take the "parent" function f(x) = x and then: 1) multiply by 5 a "stretch" 2) add -10 a "vertical translation" ? Is that what it is asking?

OpenStudy (anonymous):

No, I meant like describe how to graph this... for example: y = (x + 4)^2 − 7 Step 1 we begin with y = x^2 and replace x with x + 4 to obtain y = (x + 4)^2. The graph of y=(x+ 4)^2 is a translation of the graph of y = x^2 to the left by 4 units. like help me explain the problems above.

OpenStudy (anonymous):

helloooo....